OpenStudy (anonymous):

For which operations is the set {–1, 0, 1} closed? Choose all answers that are correct. A. addition B. division C. multiplication D. subtraction

OpenStudy (anonymous):

@chosenmatt

OpenStudy (anonymous):

I think it would be multiplaction and addition but idk

OpenStudy (chosenmatt):

i think your right

OpenStudy (anonymous):

ok i geuss i will have to go over it again in my book

OpenStudy (chosenmatt):

Solve for x. 3 • 4 – 3 • 7 = 3(x – 7) A. 4 B. –10 C. 6 D. –3

geerky42 (geerky42):

I'm sure it's only multiplication, because in addition, you can have \(1+1=2\), which is not part of set.

OpenStudy (anonymous):

ok thx geeky42

OpenStudy (anonymous):

@chosenmatt your answere is 4